# Towards the new XAI: A Hypothesis-Driven Approach to Decision Support Using Evidence

### v2

@@ 4.4. Measures
    (5) $\text{BS}_{p,i}=(C_{p,i}-A_{p,i})^{2}$
    where: $C_{p,i}$ is the likelihood level of participant $p$ in question $i$, ranging from 0 to 1; $A_{p,i}$ is the answer score of participant $p$ in question $i$, either 0 (wrong answer) or 1 (right answer). The best Brier score (i.e. equal to 0) for an individual task is when a participant answers the task correctly and gives it a 100% likelihood (or alternatively, a wrong answer but with 0% likelihoood). Therefore, a participant has better task performance when they have a lower Brier score. The Brier score measures decision accuracy, but awards a higher score for a correct answer when a participant is confident, and lower penalty for an incorrect answer when not confident. This mitigates problems where participants guess answers (i.e. have low confidence in their answers).
We then measure the over-reliance and under-reliance (Wang and Yin, 2022). Since study participants can only see the AI recommendation in C1 (Recommendation-driven), we measure agreement: whether participants have the same prediction or differ from the model’s prediction in the other two conditions.
  1. (3)
@@ 4.5. Conduct
(6) $u(h)=-\sum_{h\in H}p(h)\log p(h)$
where $u(h)$ is the uncertainty level of hypothesis $h$ given the probabilistic output is $p(h)$. To ensure there was a clear difference between high and low uncertainty, we select instances with low uncertainty by choosing instances entropy less than 0.3, and high uncertainty by choosing instances with entropy greater than 0.7. Participants did not know the certainty nor how many test instances were correct/incorrect. Each participant was paid a minimum of 4 GBP for their time, plus a bonus of 2 GBP if they could answer at least 9 out of 12 questions correctly. Participants were also given a plain language statement, and consent form and did a training phase with 3 example questions before answering the 12 test questions.
For RQ4 the text is a response to “Can you please explain why you selected this option?”. We analysed a total of 12 (questions) $\times$ 95 (participants) = 1140 responses. The final analysis includes 1,031 responses after removing 109 responses due to poor quality. Each response is assigned to at least one category (or code): Using feature values or Using evidence. We then perform a simple deductive analysis by reading each response and assign the relevant codes. We explain each code as follows: (1) Using Feature Values: participants rely on the feature values and their background knowledge to make the final decision without using the model evidence; and (2) Using Evidence: Participants rely on the evidence provided by the model and possibly their background knowledge to make the final decision.
We chose these two codes based on the idea of machine explanation and human intuition (Chen et al., 2023a, b). Specifically, using evidence refers to using the machine explanation and therefore, making use of the model evidence to support decision-making. On the other hand, using feature values is relevant to using people’s intuitions of the task based on the input feature values. Therefore, using the qualitative analysis, we explore how people use the model evidence and their intuitions in the three decision-making paradigms.
4.6. Participants
@@
5. Experiment Results
In this section, we show the results of two experiments. In the first experiment, we explore whether hypothesis-driven can improve task efficiency, task performance and reduce reliance compared to recommendation-driven and AI explanation only. In the second experiment, we understand how participants used our hypothesis-driven approach differently compared to the other two baselines.
5.1. Experiment 1: Quantitative Results
We performed a Shapiro-Wilks test to check the data normality and we found that our data was not normally distributed ($p<0.05$). Therefore, we apply non-parametric Kruskal-Wallis test. We then perform post-hoc Mann-Witney U test to do pairwise comparisons. The results are visualised in Figure 2 and Figure 2. The significant differences between two conditions are highlighted in italic red in the figures where $p<0.05$.
@@ 5.1. Experiment 1: Quantitative Results
Figure 1. Completion time. Lower is better. Means represented as dots.
Figure 2. Brier score. Lower is better. Means represented as dots.
Figure 2 shows the completion time in three conditions. There is no statistically significant difference among these three conditions ($p\approx 0.9$). We reject H1a/b. This shows that hypothesis-driven does not take more time to complete the task than recommendation-driven and AI-explanation-only.
@@
We evaluate participants’ decision-making performance by using the Brier score. A lower Brier score indicates better decision accuracy. As seen in Figure 2, participants in the hypothesis-driven condition ($M=0.267$, $SD=0.063$) have a lower Brier score than the other two approaches (C1: ($M=0.290$, $SD=0.071$), C2: ($M=0.295$, $SD=0.073$)). We accept H2a/b. Therefore, hypothesis-driven helps participants be confident when they make a correct decision, and be less confident when they make a wrong decision.
@@
Figure 3. Over-reliance. Lower is better. Means represented as dots.
Figure 4. Under-reliance. Lower is better. Means represented as dots.
In Figure 4, hypothesis-driven ($M=53.30$, $SD=22.73$) reduced over-reliance significantly compared to recommendation-driven ($M=73.86$, $SD=20.91$) ($p=1.5\times 10^{-8}$, $r=0.449$). We accept H3a. Moreover, AI-explanation-only ($M=54.21$, $SD=22.51$) also reduces over-reliance compared to the recommendation-driven approach ($p=1.6\times 10^{-8}$, $r=0.450$).
@@
In Figure 4, hypothesis-driven ($M=24.42$, $SD=18.19$) significantly reduced under-reliance compared to AI-explanation-only ($M=41.25$, $SD=27.18$) ($p=1.09\times 10^{-6}$, $r=0.387$). We accept H3b. This is not surprising because we expect that participants in the AI-explanation-only condition are the most likely to ‘under-rely’ due to being unable to compare evidence across hypotheses nor see a recommendation. Recommendation-driven ($M=17.81$, $SD=20.35$) has the least under-reliance value because participants were given recommendations.
5.2. Experiment 2: Qualitative Results
Figure 5. Frequency of using evidence to make a decision.
Figure 6. Frequency of using feature values to make a decision.
In Figure 5 and 6, we illustrate the number of times that participants used feature values and evidence to make their decisions based on the text analysis.
@@ 5.2. Experiment 2: Qualitative Results
For the AI explanation only paradigm, participants often rely on the feature values and not on the evidence explanation to make a decision. This is not surprising because participants can find it difficult to interpret the evidence without seeing the label that the evidence is referred to. We attribute this to the cognitive effort to link evidence to hypotheses, leading to participants ignoring evidence and relying on input feature values to make their decisions. This is a noteworthy limitation of AI-explanation-only as it makes people overlook the explanation if the link to the evidence is unclear. In the study by Gajos and Mamykina (2022), the link from feature attributions to the task solution is more straightforward than in our study, which may explain the divergence of results.
Participants more often use the evidence to make a decision in the hypothesis-driven than in recommendation-driven or AI explanation only. This shows participants took advantage of the model evidence. In the two baseline conditions, participants tended to ignore evidence seemingly due to the inability to interpret it, which means they will fail to take advantage of the underlying model. In Figure 5, there are only two exceptions at Q6 and Q11 where the evidence is not the most used in the hypothesis-driven condition.
Figure 7. An example of uncertainty awareness (Q6).
Figure 8. An example of deceptive evidence (Q9).
@@
In summary, the qualitative analysis showed that participants took advantage of the decision aid more in the hypothesis-driven condition than in recommendation-driven and explanation-only conditions. Further, we also found that participants recognised model uncertainty in the hypothesis-driven condition. However, there still remains a limit of having deceptive evidence.
6. Discussions and Conclusions
@@
Moreover, hypothesis-driven reduced over-reliance significantly compared to the standard AI recommendation. Similarly, hypothesis-driven also reduced under-reliance compared to AI explanation only. Importantly, the positive result for under-reliance using recommendation-driven is not cancelled out by the poor over-reliance result, compared to hypothesis-driven. The primary aim indicating potential for the use of uncertainty/confidence (Bhatt et al., 2021; Le et al., 2023) and conformal prediction (Straitouri et al., 2023) to direct decision makers’ attention towards a set of hypotheses that it is confident about.
@@
There are also some limitations with the study. First, we ran the experiment on one dataset (Ames Housing), which limits generalisability. In addition, as there is no ground truth for the price of a house, the experimental participants’ tasks are somewhat subjective. Further, this task has only three output classes, so only three hypotheses, and we anticipate the results would be more interesting when we consider more hypotheses. Finally, the human experiment is currently conducted with laypeople while experts likely interact with the decision-aiding tool differently from laypeople (Fogliato et al., 2022).
7. Conclusions
@@ Appendix A Statistics of Experiment 1
(C1) Recommendation-driven 102.000 1075.186 560.534 306.000 689.500 946.000 1307.500 4021.000
(C2) AI Explanation Only 99.000 1117.162 675.505 210.000 742.500 928.000 1345.000 5559.000
(C3) Hypothesis-driven 101.000 1085.228 555.319 358.000 711.000 944.000 1287.000 3350.000
Table 1. Statistics of completion time per condition.
@@
Table 2. Statistics of Brier score per condition.
@@
Table 3. Statistics of over-reliance per condition.
@@
Table 4. Statistics of under-reliance per condition.
@@ Appendix B Subjective Questions in Experiment 1
We evaluate Q1-4 separately to measure 4 measures (In control, Preference, Mental demand and System complexity). We aggregate Q5-12 to measure Trust. We based the trust questions on questions from (Hoffman et al., 2018).
In Figure 9, AI-explanation-only is significantly worse than the other conditions in all facets. Moreover, recommendation-driven and hypothesis-driven are quite similar and there is no statistical difference between these two conditions. The reason is that we conduct between-subject experiments so each participant has access to only one condition and they do not have another condition to compare to. If we run within-subject experiments to measure subjective questions in the future, participants can compare different decision-making approaches and evaluate which one they prefer the most.
(a) In control (Higher is better)
(b) Preference (Higher is better)
(c) Mental demand (Lower is better)
(d) System complexity (Lower is better)
(e) Trust (Higher is better)
Figure 9. Subjective Measures in Experiment 1. Means represented as dots.
@@ Appendix C Example Questions in Human Experiment
First, Figure 10(a)11(a) and 12(a) show the first page of the training phase in three conditions. This page introduces the task, how to read the evidence, what information will be given and how the participants will be scored. After that, three example questions will be shown, one of which is Figure 10(b) or 11(b) or 12(b) depending on the condition.
(a) Introduction page
(b) An example question
Figure 10. Training phase in (C1) Recommendation-driven
(a) Introduction page
(b) An example question
Figure 11. Training phase in (C2) AI-explanation-only
(a) Introduction page
(b) An example question
Figure 12. Training phase in (C3) Hypothesis-driven
@@
Example questions in Experiment 1 are shown in Figure 13, 14 and 15. Experiment 2 will use the same set of questions, with an addition of having a free text after the likelihood slider to ask participants why they made such choices.
@@
  • In Figure 15, the house features selected are similar to the example question in Figure 13 and 14. We show participants the evidence for all hypotheses (low, medium and high). We do not give them the AI prediction. Specifically, we have supportive evidence in most features for hypothesis low. By contrast, strongly negative evidence refute hypothesis high. The correct answer here is low.
Figure 13. A screenshot of a question in (C1) Recommendation-driven.
Figure 14. A screenshot of a question in (C2) AI-explanation-only.
Figure 15. Screenshots of evidence provided for all three hypotheses in (C3) Hypothesis-driven.

### v3

@@ 4.4 Measures
    $\text{BS}_{p}=\frac{1}{N}\sum_{i=1}^{N}(C_{p,i}-A_{p,i})^{2}$ (5)
    where: $C_{p,i}$ is the likelihood level of participant $p$ in question $i$, ranging from 0 to 1; $A_{p,i}$ is the answer score of participant $p$ in question $i$, either 0 (wrong answer) or 1 (right answer); N is the number of questions for each participant. The best Brier score (i.e. equal to 0) for an individual task is when a participant answers the task correctly and gives it a 100% likelihood (or alternatively, a wrong answer but with 0% likelihoood). Therefore, a participant has better task performance when they have a lower Brier score. The Brier score measures decision accuracy, but awards a higher score for a correct answer when a participant is confident, and lower penalty for an incorrect answer when not confident. This mitigates problems where participants guess answers (i.e. have low confidence in their answers).
We then measure the over-reliance and under-reliance [39]. Since study participants can only see the AI recommendation in C1 (Recommendation-driven), we measure agreement: whether participants have the same prediction or differ from the model’s prediction in the other two conditions.
  1. 3.
@@ 4.5 Conduct
$u(h)=-\sum_{h\in H}p(h)\log p(h)$ (6)
where $u(h)$ is the uncertainty level of hypothesis $h$ given the probabilistic output is $p(h)$. To ensure there was a clear difference between high and low uncertainty, we select instances with low uncertainty by choosing instances entropy less than 0.3, and high uncertainty by choosing instances with entropy greater than 0.7. Participants did not know the certainty nor how many test instances were correct/incorrect. Each participant was paid a minimum of £4 for their time, plus a bonus of £2 if they could answer at least 9 out of 12 questions correctly. Participants were also given a plain language statement, and consent form and did a training phase with 3 example questions before answering the 12 test questions.
For RQ4 the text is a response to “Can you please explain why you selected this option?". We analysed a total of 12 (questions) $\times$ 95 (participants) = 1140 responses. The final analysis includes 1,031 responses after removing 109 responses due to poor quality. Each response is assigned to at least one category (or code): Using feature values or Using evidence. We then perform a simple deductive analysis by reading each response and assign the relevant codes. We explain each code as follows: (1) Using Feature Values: participants rely on the feature values and their background knowledge to make the final decision without using the model evidence; and (2) Using Evidence: Participants rely on the evidence provided by the model and possibly their background knowledge to make the final decision.
We chose these two codes based on the idea of machine explanation and human intuition [8, 9]. Specifically, using evidence refers to using the machine explanation and therefore, making use of the model evidence to support decision-making. On the other hand, using feature values is relevant to using people’s intuitions of the task based on the input feature values. Therefore, using the qualitative analysis, we explore how people use the model evidence and their intuitions in the three decision-making paradigms.
4.6 Participants
@@
5 Experiment Results
In this section, we show the results of two experiments. In the first experiment, we explore whether hypothesis-driven can improve task efficiency, task performance and reduce reliance compared to recommendation-driven and AI explanation only. In the second experiment, we understand how participants used our hypothesis-driven approach differently compared to the other two baselines. The full statistics and screenshots from our user study can be found in the supplementary material [25].
R O H R vs. O R vs. H O vs. H
Time (minutes) $\mathbf{17.92}$  $\pm$ $\mathbf{9.34}$ 18.62  $\pm$ 11.26 18.09  $\pm$ 9.26 $p=0.808,r=0.020$ $p=0.993,r=0.001$ $p=0.892,r=0.011$
Brier score 0.29  $\pm$ 0.07 0.30  $\pm$ 0.07 $\mathbf{0.27}$  $\pm$ $\mathbf{0.06}$ $p=0.566,r=0.047$ $\mathbf{p=0.016,r=0.196}$ $\mathbf{p=0.003,r=0.245}$
Over-reliance (%) 73.86  $\pm$ 20.91 54.21  $\pm$ 22.51 $\mathbf{53.30}$  $\pm$ $\mathbf{22.73}$ $\mathbf{p<0.001,r=0.450}$ $\mathbf{p<0.001,r=0.449}$ $p=0.946,r=0.006$
Under-reliance (%) $\mathbf{17.81}$  $\pm$ $\mathbf{20.35}$ 41.25  $\pm$ 27.18 24.42  $\pm$ 18.19 $\mathbf{p<0.001,r=0.523}$ $\mathbf{p=0.001,r=0.246}$ $\mathbf{p<0.001,r=0.387}$
Table 1: Results of the human experiment. R: Recommendation-driven, O: AI-explanation-only, H: Hypothesis-driven. Winners/significances are highlighted.
5.1 Experiment 1: Quantitative Results
We performed a Shapiro-Wilks test to check the data normality and we found that our data was not normally distributed ($p<0.05$). Therefore, we apply non-parametric Kruskal-Wallis test. We then perform post-hoc Mann-Witney U test to do pairwise comparisons. The results are visualised in Table 1 and Table 1. The significant differences between two conditions are highlighted in italic red in the figures where $p<0.05$.
@@ 5.1 Experiment 1: Quantitative Results
Table 1 shows the completion time in three conditions. There is no statistically significant difference among these three conditions ($p\approx 0.9$). We reject H1a/b. This shows that hypothesis-driven does not take more time to complete the task than recommendation-driven and AI-explanation-only.
@@
We evaluate participants’ decision-making performance by using the Brier score. A lower Brier score indicates better decision accuracy. As seen in Table 1, participants in the hypothesis-driven condition ($M=0.267$, $SD=0.063$) have a lower Brier score than the other two approaches (C1: ($M=0.290$, $SD=0.071$), C2: ($M=0.295$, $SD=0.073$)). We accept H2a/b. Therefore, hypothesis-driven helps participants be confident when they make a correct decision, and be less confident when they make a wrong decision.
@@
In Table 1, hypothesis-driven ($M=53.30$, $SD=22.73$) reduced over-reliance significantly compared to recommendation-driven ($M=73.86$, $SD=20.91$) ($p=1.5\times 10^{-8}$, $r=0.449$). We accept H3a. Moreover, AI-explanation-only ($M=54.21$, $SD=22.51$) also reduces over-reliance compared to the recommendation-driven approach ($p=1.6\times 10^{-8}$, $r=0.450$).
@@
In Table 1, hypothesis-driven ($M=24.42$, $SD=18.19$) significantly reduced under-reliance compared to AI-explanation-only ($M=41.25$, $SD=27.18$) ($p=1.09\times 10^{-6}$, $r=0.387$). We accept H3b. This is not surprising because we expect that participants in the AI-explanation-only condition are the most likely to ‘under-rely’ due to being unable to compare evidence across hypotheses nor see a recommendation. Recommendation-driven ($M=17.81$, $SD=20.35$) has the least under-reliance value because participants were given recommendations.
5.2 Experiment 2: Qualitative Results
Figure 1: Frequency of using evidence to make a decision.
Figure 2: Frequency of using feature values to make a decision.
In Figure 1 and 2, we illustrate the number of times that participants used feature values and evidence to make their decisions based on the text analysis.
@@ 5.2 Experiment 2: Qualitative Results
For the AI explanation only paradigm, participants often rely on the feature values and not on the evidence explanation to make a decision. This is not surprising because participants can find it difficult to interpret the evidence without seeing the label that the evidence is referred to. We attribute this to the cognitive effort to link evidence to hypotheses, leading to participants ignoring evidence and relying on input feature values to make their decisions. This is a noteworthy limitation of AI-explanation-only as it makes people overlook the explanation if the link to the evidence is unclear. In the study by Gajos and Mamykina [15], the link from feature attributions to the task solution is more straightforward than in our study, which may explain the divergence of results.
Participants more often use the evidence to make a decision in the hypothesis-driven approach than in recommendation-driven or AI explanation only. This shows participants took advantage of the model evidence. In the two baseline conditions, participants tended to ignore evidence seemingly due to the inability to interpret it, which means they will fail to take advantage of the underlying model. In Figure 1, there are only two exceptions at Q6 and Q11 where the evidence is not the most used in the hypothesis-driven condition.
Figure 3: An example of uncertainty awareness (Q6).
Figure 4: An example of deceptive evidence (Q9).
@@
In summary, the qualitative analysis showed that participants took advantage of the decision aid more in the hypothesis-driven condition than in recommendation-driven and explanation-only conditions. Further, we also found that participants recognised model uncertainty in the hypothesis-driven condition. However, there still remains a limitation of having deceptive evidence.
6 Discussions
@@
Moreover, hypothesis-driven reduced over-reliance significantly compared to the standard AI recommendation. Similarly, hypothesis-driven also reduced under-reliance compared to AI explanation only. Importantly, the positive result for under-reliance using recommendation-driven is not cancelled out by the poor over-reliance result, compared to hypothesis-driven. The primary aim indicating potential for the use of uncertainty/confidence [5, 24] and conformal prediction [33] to direct decision makers’ attention towards a set of hypotheses that it is confident about.
@@
There are also some limitations with the study. First, we ran the experiment on one dataset (Ames Housing), which limits generalisability. In addition, as there is no ground truth for the price of a house, the experimental participants’ tasks are somewhat subjective. Further, this task has only three output classes, so only three hypotheses, and we anticipate the results would be more interesting when we consider more hypotheses. Finally, the human experiment is currently conducted with laypeople while experts likely interact with the decision-aiding tool differently from laypeople [14].
7 Conclusions
@@ Appendix A Statistics of Experiment 1
(C1) Recommendation-driven 102.000 17.920 9.342 5.100 11.492 15.767 21.792 67.017
(C2) AI-explanation-only 99.000 18.619 11.258 3.500 12.375 15.467 22.417 92.650
(C3) Hypothesis-driven 101.000 18.087 9.255 5.967 11.850 15.733 21.450 55.833
Table 2: Statistics of completion time per condition (in minutes).
@@
Table 3: Statistics of Brier score per condition.
@@
Table 4: Statistics of over-reliance (%) per condition.
@@
Table 5: Statistics of under-reliance (%) per condition.
In this section, we show the statistics of four measures in Experiment 1 as in Table 2,  3,  4 and  5. We also show the boxplot of these measures in Figure 5,  6,  7 and  8.
Figure 5: Completion time. Lower is better. Means represented as dots.
Figure 6: Brier score. Lower is better. Means represented as dots.
Figure 7: Over-reliance. Lower is better. Means represented as dots.
Figure 8: Under-reliance. Lower is better. Means represented as dots.
@@ Appendix B Subjective Questions in Experiment 1
We evaluate Q1-4 separately to measure 4 measures (In control, Preference, Mental demand and System complexity). We aggregate Q5-12 to measure Trust. We based the trust questions on questions from [18].
Based on Figure 9,  10,  11,  12 and  13, AI-explanation-only is significantly worse than the other conditions in all facets. Moreover, recommendation-driven and hypothesis-driven are quite similar and there is no statistical difference between these two conditions. The reason is that we conduct between-subject experiments so each participant has access to only one condition and they do not have another condition to compare to. If we run within-subject experiments to measure subjective questions in the future, participants can compare different decision-making approaches and evaluate which one they prefer the most.
Figure 9: In control (Higher is better)
Figure 10: Preference (Higher is better)
Figure 11: Mental demand (Lower is better)
Figure 12: System complexity (Lower is better)
Figure 13: Trust (Higher is better)
@@ Appendix C Example Questions in Human Experiment
First, Figure 1416 and 18 show the first page of the training phase in three conditions. This page introduces the task, how to read the evidence, what information will be given and how the participants will be scored. After that, three example questions will be shown, one of which is Figure 15 or 17 or 19 depending on the condition.
Figure 14: Introduction page in (C1) Recommendation-driven
Figure 15: Training phase in (C1) Recommendation-driven: An example question
Figure 16: Introduction page in (C2) AI-explanation-only
Figure 17: Training phase in (C2) AI-explanation-only: An example question
Figure 18: Introduction page in (C3) Hypothesis-driven
Figure 19: Training phase in (C3) Hypothesis-driven: An example question
@@
Example questions in Experiment 1 are shown in Figure 20, 21 and 22. Experiment 2 will use the same set of questions, with an addition of having a free text after the likelihood slider to ask participants why they made such choices.
@@
  • In Figure 22, the house features selected are similar to the example question in Figure 20 and 21. We show participants the evidence for all hypotheses (low, medium and high). We do not give them the AI prediction. Specifically, we have supportive evidence in most features for hypothesis low. By contrast, strongly negative evidence refute hypothesis high. The correct answer here is low.
Figure 20: A screenshot of a question in (C1) Recommendation-driven.
Figure 21: A screenshot of a question in (C2) AI-explanation-only.
Figure 22: Screenshots of evidence provided for all three hypotheses in (C3) Hypothesis-driven.